OpenStudy (fuzzyball):

What is the point-slope form of the equation of a line that passes through the point (3, 5) and has a slope of 2?

OpenStudy (3mar):

May I help?

OpenStudy (fuzzyball):

yea

OpenStudy (3mar):

Thanks

OpenStudy (3mar):

The form is \[\frac{ y_2-y_1 }{ x_2-x_1 }=\frac{ y-y_1 }{ x-x_1 }=m\] where m is the slope and \[(x_1,y_1)\] is the point that the line passes through

OpenStudy (3mar):

so you can now substitute with x1 and y1 and m u=into that equation, and let me know what you got

OpenStudy (fuzzyball):

ok

OpenStudy (fuzzyball):

give me a min or two im slow sorry

OpenStudy (3mar):

m u=into *m into

OpenStudy (3mar):

No problem .. Take your time.

OpenStudy (fuzzyball):

2= y-5/x-3

OpenStudy (fuzzyball):

I WILL Medal and fan..if you help me

OpenStudy (3mar):

Great!

OpenStudy (3mar):

\[y-5=2(x-3)\] \[y-5=2x-6\] \[y=2x-1\]

OpenStudy (3mar):

Medals don't matter. What matters is that got the idea and can solve such this yourself.